OpenStudy (anonymous):

If y varies directly as x and y = 14 when x = 8, find x when y = 21. A. 24 B. 12 C. 36 D. 16

OpenStudy (anonymous):

y varies directly as x: y ~ x or can be also written as y = kx where k is a constant scalar value given: y = 14 when x = 8 y = kx or k = y/x = 14/8 = 7/4 k = 7/4 find x when y = 21 y = k*x x = y/k = 21/(7/4) = 21 *4 /7 = ?

OpenStudy (anonymous):

@drewsorr1 that is actually really confusing :p

OpenStudy (anonymous):

Hm, then I'll let mathmale help you understand.

OpenStudy (anonymous):

@drewsorr1 sorry

OpenStudy (mathmale):

There are two general formulas that you should know for direct and inverse variation. In the case of direct variation, the dependent variable (usually y) increases when the independent variable (usually x) increases. That's what we have here. In the case of inverse variation, y decreases when x increases. OK. You're told that "y varies directly as x." This is the same thing as "direct variation." The proper formula for this is y=kx, where k is some constant, which we call the "proportionality constant."

OpenStudy (mathmale):

Start with this model. y = k*x. Your posted math problem states that " y = 14 when x = 8." Let y=14 and x=8 in this model. Can you solve the resulting equation for k?

OpenStudy (anonymous):

is this correct?